OpenStudy (anonymous):

Hello. What would my final grade be if my current grade is 86% and fail a final which is worth 20% of my grade? Thank you!

OpenStudy (studio-games):

You shouldnt use that as an excuse to fail a test though.

OpenStudy (studio-games):

Try your hardest @Mheyam !

OpenStudy (anonymous):

i plan to give it my all. i just wanted to ask just incase my luck runs out that day. sad times

OpenStudy (studio-games):

I understand. I hope you get a great grade on it! Good luck with it!

OpenStudy (studio-games):

Remember that openstudy will be here for you if you need help on your test!

OpenStudy (fortytherapper):

What's your target grade?

jigglypuff314 (jigglypuff314):

I messed up before, sorry it is 86% is only (100-20)% of your grade at the moment Final grade = Grade times weight Final grade = (0.86 * 80) + (grade * 20) so for example, if you get a 53% on your final your final grade would be = (0.86 * 80) + (0.53 * 20)

OpenStudy (mjdennis):

Stealing @jigglypuff314 's math: Final grade = (86 * 0.80) + (grade * 0.20) Turn it around to get your targets: test_target_grade = 5*(final_target_grade-68.8) If you can put the formula in a spreadsheet, it tells you all sorts of interesting stuff. What I find most interesting is that you need 106 on the exam to get 90 for the course (so an A looks out of the question), but B+ is still achievable! Do your best!